OpenStudy (anonymous):

Find the first four terms of the sequence given by the following. An = 3^n + 4 /3^n N = 1,2,3,...

OpenStudy (anonymous):

A1 = 3 + 4/3 A2 = 3^2 + 4/3^2 = 9 + 4/9 A3 = 3^3 + 4/3^3 = 27 + 4/27 A4 =

OpenStudy (anonymous):

3^{n}+ \dfrac{4}{3^{n}}

OpenStudy (anonymous):

So these would be my answers?

OpenStudy (anonymous):

Once you have used the info to figure out A4.

OpenStudy (anonymous):

Okay, Thank you! (:

OpenStudy (anonymous):

You are very welcome. Glad to have helped.